# Generative AI for Game Theory-based Mobile Networking

### v1

@@ V-A GDM-enhanced Non-Cooperative Game
For non-cooperative game scenarios in networking with numerous players and complex decision spaces, obtaining a Nash equilibrium is often challenging computationally due to the need to search through a high-dimensional, exponentially large space. GDM has outstanding capabilities in modeling complex data distribution and data generation [14]. With the help of GDM, it is possible to enhance the search efficiency for Nash equilibrium and greatly reduce the computational complexity of complex non-cooperative games.
@@ V-B Performance Evaluation of GAI-enabled Game Theory
It is crucial to design a comprehensive performance evaluation model to evaluate the performance of GAI combined with game theory. This is because effective evaluation models can provide directions for guiding and improving the GAI-enabled game theory framework. Furthermore, intelligent evaluation criteria should be designed to evaluate GAI-enabled game theory instead of human evaluation in the future.
@@ V-C GAI-enabled Game Theory for Future Networks
As a key technology for future 6G wireless communications, the space-air-ground integrated network (SAGIN) has been proposed to provide seamless communication coverage and high-speed connections [15]. Due to the inclusion of heterogeneous network technologies, i.e., ground-based network, air-based network, and space-based network, the deployment and network management of SAGIN becomes more challenging. The vast potential of GAI-enabled game theory in networking is anticipated to synergize with the emerging SAGIN technology, catalyzing the advancement of SAGIN.
@@ VI Conclusion
In this article, we explored the integration of LLMs with game theory, and concluded the advantages of LLM for game theory. Following this, we presented the potential applications of LLM combined with game theory in networking. Based on these, we proposed a framework that utilized RAG to combine LLM with game theory, aiming to achieve the intelligence of game theory in network applications. We conducted a case study on UAV secure communication optimization to validate the effectiveness of the proposed framework. Finally, we discussed several potential research directions for future extensions.

### v2

@@ V-A GDM-enhanced Non-Cooperative Game
For the non-cooperative game for the networking scenarios with numerous players and complex decision spaces, obtaining a Nash equialibrium is often computationally challenging due to the need to search through a high-dimensional and exponentially large space. GDM has outstanding capabilities in modeling complex data distribution and data generation [14]. With the help of GDM, it is possible to enhance the search efficiency for Nash equilibrium. Therefore, future work could focus on incorporating GDM into the solution process of non-cooperative games, aiming to identify solutions that satisfy Nash equilibrium conditions through sampling and optimization of the strategy space.
@@ V-B Performance Evaluation of GAI-enabled Game Theory
It is crucial to design a comprehensive performance evaluation model to evaluate the performance of GAI combined with game theory. This is because effective evaluation models can provide directions for guiding and improving the GAI-enabled game theory framework. However, traditional approaches often rely on human expertise to formulate evaluation models, which tend to lack flexibility. Therefore, future work could explore the integration of human feedback with existing evaluation models by utilizing LLMs and RAG technology to design more intelligent evaluation models.
@@ V-C GAI-enabled Game Theory for Future Networks
As a key technology for future 6G wireless communications, the space-air-ground integrated network (SAGIN) has been proposed to provide seamless communication coverage and high-speed connections [15]. However, due to the heterogeneous network components and the dynamic network environment, optimizing SAGIN is a challenging task. Therefore, future research could explore the use of GAI-enabled game theory to model and optimize network scheduling, routing strategies, and resource allocation within the dynamic network environment of SAGIN.
@@ VI Conclusion
In this article, we explored the integration of LLMs with game theory, and concluded the advantages of LLMs for game theory. Following this, we presented the potential applications of LLMs combined with game theory in networking. Based on these, we proposed a framework that utilized RAG to combine LLMs with game theory, aiming to achieve the intelligence of game theory in network applications. We conducted a case study on UAV secure communication optimization to validate the effectiveness of the proposed framework. Finally, we discussed several potential research directions for future extensions.
Biographies
Long He (helong0517@foxmail.com) received a BS degree in Computer Science and Technology from Chengdu University of Technology, Sichuan, China, in 2019. He is currently working toward the PhD degree in Computer Science and Technology at Jilin University, Changchun, China. His research interests include vehicular networks and edge computing.
Geng Sun [S’17, M’19, SM’24] (sungeng@jlu.edu.cn) received a B.S. degree in Communication Engineering from Dalian Polytechnic University, China, and the Ph.D. degree in Computer Science from Jilin University, China, in 2011 and 2018, respectively. He was a visiting researcher in the School of Electrical and Computer Engineering at Georgia Institute of Technology, USA. He is currently a Professor in the College of Computer Science and Technology at Jilin University, and his research interests include UAV Networks, collaborative beamforming, generative AI and optimizations.
Dusit Niyato [M’09, SM’15, F’17] (dniyato@ntu.edu.sg) is currently a professor in the School of Computer Science and Engineering, Nanyang Technological University, Singapore. He received the B.Eng. degree from King Mongkut’s Institute of Technology Ladkrabang (KMITL), Thailand, in 1999, and the Ph.D. in electrical and computer engineering from the University of Manitoba, Canada, in 2008. His research interests are in the areas of Internet of Things (IoT), machine learning, and incentive mechanism design.
Hongyang Du (duhy@eee.hku.hk) is an assistant professor at the Department of Electrical and Electronic Engineering, the University of Hong Kong. He received his Ph.D. degree from the College of Computing and Data Science, Energy Research Institute @ NTU, Nanyang Technological University, Singapore, in 2024. He serves as the Editor-in-Chief assistant of IEEE Communications Surveys & Tutorials (2022-2024), and the Guest Editor for IEEE Vehicular Technology Magazine. His research interests include edge intelligence, generative AI, and network management.
Fang Mei (meifang@jlu.edu.cn) received the M.Sc. and Ph.D. degrees in computer science from Jilin University, Changchun, China, in 2005 and 2010, respectively. She is currently an Associate Professor in the College of Computer Science and Technology at Jilin University. Her research interests include intelligent information processing, multi-access edge computing, vehicular networks, and antenna arrays.
Jiawen Kang (kavinkang@gdut.edu.cn) received the Ph.D. degree from the Guangdong University of Technology, China, in 2018. He has been a postdoc at Nanyang Technological University, Singapore, from 2018 to 2021. He is currently a full professor at Guangdong University of Technology, China. His research interests focus on blockchain, security, and privacy protection.
Mérouane Debbah [S’01, M’04, SM’08, F’15] (merouane.debbah@ku.ac.ae) received the M.Sc. and Ph.D. degrees from the Ecole Normale Supérieure Paris-Saclay, France. Since 2023, he is a Professor at Khalifa University of Science and Technology in Abu Dhabi and founding director of the 6G center. He has managed 8 EU projects and more than 24 national and international projects. His research interests lie in fundamental mathematics, algorithms, statistics, information, and communication sciences research. He holds more than 50 patents. He is an IEEE Fellow, a WWRF Fellow, a Eurasip Fellow, an AAIA Fellow, an Institut Louis Bachelier Fellow and a Membre émérite SEE.
Zhu Han [S’01, M’04, SM’09, F’14] (zhuhan22@gmail.com) currently is a professor in the Electrical and Computer Engineering Department at the University of Houston, Texas. He has been an AAAS Fellow since 2019. He received the IEEE Kiyo Tomiyasu Award in 2020. He has been a 1 percent highly cited researcher since 2017 according to Web of Science.